# Soft Reasoning on Uncertain Knowledge Graphs

### v1

@@ Abstract
The study of machine learning-based logical query-answering enables reasoning with large-scale and incomplete knowledge graphs. This paper further advances this line of research by considering the uncertainty in the knowledge. The uncertain nature of knowledge is widely observed in the real world, but does not align seamlessly with the first-order logic underpinning existing studies. To bridge this gap, we study the setting of soft queries on uncertain knowledge, which is motivated by the establishment of soft constraint programming. We further propose an ML-based approach with both forward inference and backward calibration to answer soft queries on large-scale, incomplete, and uncertain knowledge graphs. Theoretical discussions present that our methods share the same complexity as state-of-the-art inference algorithms for first-order queries. Empirical results justify the superior performance of our approach against previous ML-based methods with number embedding extensions.
complex query answering, uncertain knowledge graph, neural link predictor
@@ 1 Introduction
Representing and reasoning with factual knowledge are essential functionalities of artificial intelligence systems. As a powerful way of knowledge representation, knowledge graphs (KGs) (Miller, 1995; Suchanek et al., 2007; Vrandečić & Krötzsch, 2014) use nodes to represent entities and edges to encode the relations between entities. The reasoning process on KGs can be formally justified by answering logic queries. The key challenge is the open world assumption, where the KG is assumed to be incomplete (Libkin & Sirangelo, 2009). Thus, the approaches for the closed world data, such as graph matching methods or those for relational database queries, are prohibited. In recent studies on logical queries on knowledge graphs, the generalizability of machine learning models is leveraged to predict the missing information in the open world and conduct first-order logic reasoning (Ren & Leskovec, 2020; Arakelyan et al., 2021; Liu et al., 2021; Wang et al., 2023b). This combination of machine learning and logic enables further possibilities in data management (Wang et al., 2022; Ren et al., 2023).
Uncertain knowledge is widely observed from the daily events (Zhang et al., 2020) to the interaction of biological systems (Szklarczyk et al., 2023). Besides, uncertainty is also particularly pervasive in KGs because KGs are constructed by information extraction models that could introduce errors (Angeli et al., 2015; Ponte & Croft, 2017) and from large corpses that could be noisy (Carlson et al., 2010). To represent the uncertain knowledge, confidence values $p$ are associated with triples in many well-established KGs (Carlson et al., 2010; Speer et al., 2017; Szklarczyk et al., 2023), known as the uncertain KG. Notably, recent studies complete the missing triples and their uncertain values with the generalization power of ML models (Chen et al., 2019; Pai & Costabello, 2021).
To reason with uncertainty, many extensions of first-order logic have been made to cope with the uncertainty in knowledge representation systems formally (Adams, 1996). It is noteworthy to mention that the study of probabilistic databases also extends the relational databases with a confidence value $p\in[0,1]$ (Cavallo & Pittarelli, 1987; Suciu et al., 2022). From a machine learning perspective, however, previous studies in open world probabilistic databases are limited from two aspects: (1) They assume uniform uncertainty for all unobserved knowledge (Ceylan et al., 2021), leading to weaker characterization of the open world knowledge without the generalization (2) They focus on first-order logic, which might be insufficient to describe practical reasoning processes with uncertainty, as shown in Example 1.
@@
Figure 1: (a) Examples of two soft queries in the candidate search procedure. The soft queries introduced in this paper are jointly defined by first-order logic and soft requirements. Particularly, soft requirements (necessity and importance) are introduced to characterize fine-grained decision-making preferences, distinguishing them from first-order queries. (b) Open world uncertain KG for to what extent a candidate possesses a skill. Solid lines indicate the observed knowledge, while dashed lines indicate the unobserved data. Values indicate uncertainty, whereas the higher value indicates the fact is more likely to be true.
Example 1 (Soft query models candidate search criteria).
Figure 1 demonstrates the soft queries to model the employers’ preferences for two roles of jobs. The $[0,1]$ values in uncertain KG describe how well a candidate masters a specific skill. We can see that the two jobs have different criteria. The necessity from $[0,1]$ suggests the minimum requirement for the condition to be satisfied. The importance from $[0,\infty)$ indicates the importance of a constraint.
This paper studies the machine learning approach for reasoning with open world and uncertain knowledge, which advances previous studies in symbolic probabilistic databases. Its contribution is threefold.
@@
Problem Settings Language of reasoning Uncertainty of Knowledge Open World Assumption (Unobserved Knowledge)
@@
Logical queries on KG EFO - ML generalization to unobserved facts
@@
Contribution 1: A novel and practical setting. We propose a novel setting of Soft Queries on Uncertain KG (SQUK). Our setting extends the previous setting of logical queries on KGs in two ways: (1) For the underlying knowledge base, SQUK extends the open world KG to open world and uncertain KG. (2) For the language describing reasoning, SQUK extends first-order language to uncertainty-aware soft queries with soft requirements, which are motivated by real-world reasoning with uncertainty (see Example 1) and the establishment in soft constraint programming (Schiex, 1992; Rossi et al., 2006). The formal definition will be introduced in Section 3. The comparison of SQUK against other settings is detailed in Table 1.
@@
Contribution 2: ML method for soft queries. We bridge machine learning and SQUK by proposing Soft Reasoning with calibrated Confidence values (SRC), which uses Uncertain Knowledge Graph Embeddings (UKGEs) to tackle the open world assumption and achieves the same computational complexity as the state-of-the-art inference algorithms (Bai et al., 2023; Yin et al., 2024). The error analysis is also conducted for SRC under mild assumptions, characterizing how the performance is affected by UKGEs and the query structures. Based on our analysis, we suggest calibrating the confidence by debiasing and learning, which further boosts the performance of SRC.
@@
Contribution 3: Extensive empirical studies. We also conduct extensive empirical studies to benchmark the performance of a broad spectrum of methods under the UKGE (Chen et al., 2019) settings. SRC is compared against baselines including Query Embedding (QE) (Ren & Leskovec, 2020) methods with Number Embeddings (QE+NE) (Vaswani et al., 2017). In particular, we compared the differences between QE+NE and SRC under various soft query settings, demonstrating the advantage of SRC. We also make a fair comparison with large language models on annotated soft queries in natural language setting.
We present the related work in Appendex A.
2 Preliminary
This section presents the problem setting of First-Order (FO) query answering on open world knowledge graphs, which previous studies have widely investigated. Then, we briefly discuss two ML-based approaches: query embeddings and inference methods. We note that existing studies are not compatible with the uncertainty in knowledge.
2.1 Logical queries on knowledge graphs
Definition 1 (Knowledge graphs).
Let $\mathcal{E}$ be the set of entities and $\mathcal{R}$ be the set of relations. A knowledge graph is a set of triples $\mathcal{KG}=\{(s_{i},r_{i},o_{i})\}$, where $s_{i},o_{i}\in\mathcal{E}$ are entities and $r\in\mathcal{R}$ is relation.
The fundamental challenge for knowledge graphs is the Open World Assumption (OWA), which assumes that the observed KG $\mathcal{KG}$ is a subset of the underlying KG $\mathcal{KG}_{\rm OWA}$.
Definition 2 (Open world assumption).
$\mathcal{KG}\subset\mathcal{KG}_{\rm OWA}$.
The study of logical queries on KG focuses on Existential First-Order (EFO) queries, usually with one free variable.
Definition 3 (Syntax of existential first-order queries).
Let $\Gamma$ be an existential first-order query in the Disjunctive Normal Form (DNF) (Marker, 2006):
$\Gamma(y)=\gamma_{1}(y)\lor\cdots\lor\gamma_{q}(y),$ (1)
The notation $\Gamma(s/y)$ means we assign the answer entity $s$ to the variable $y$. When all free variables are substituted, a query $\Gamma(y)$ is transformed into a sentence $\Gamma(s/y)$. Given $\mathcal{KG}$, answering a query $\Gamma(y)$ means finding all substitutions, such that the sentence $\Gamma(s/y)$ is entailed by $\mathcal{KG}$, i.e., $\mathcal{KG}\models\Gamma(s/y)$. The answer set is defined as
Definition 4 (Answer set of first-order query).
Given $\mathcal{KG}$ and a query $\Gamma(y)$, the answer set is
$A=\{s:\mathcal{KG}\models\Gamma(s/y)\}.$ (3)
The entailment symbol $\models$ is conducted under the standard calculus of logical connectives with boolean semantics. In short, check $\models$ operation is broken down to check whether each substituted atomic query is a triple in $\mathcal{KG}$. For a detailed definition, we refer readers to Marker (2006).
Challenge of Open World Assumption. Similarly, the answer set in the open world is
$A_{\rm OWA}=\{s:\mathcal{KG}_{\rm OWA}\models\Gamma(s/y)\},$ (4)
which is different from $A$ in the closed world. Therefore, traditional graph-matching methods fail because of the generalizability (Riesen et al., 2010).
2.2 ML-based method for logical queries on KG
Recent works are dedicated to introducing ML methods, i.e., knowledge graph embeddings, to generalize from $\mathcal{KG}$ to $\widehat{\mathcal{KG}}$, so that it approximates $\mathcal{KG}_{\rm OWA}$.
Query Embeddings (QE). Query embedding methods generally map the query $\Gamma$ as an embedding (Ren & Leskovec, 2020; Liu et al., 2021; Wang et al., 2023b). One dominant way is to “translate” the procedure of solving logical formulas in Equation (1) into the set operations, such as set projection, intersection, union, and complement. Then, the neural networks are designed to model such set operations in the embedding space. We can see that direct modeling of set operations is incompatible with the concept of necessity and importance introduced in Section 3.
Inference methods. Other methods solve the open world query answering methods in a two-step approach (Arakelyan et al., 2021; Bai et al., 2023; Yin et al., 2024). In the first step, the pre-trained knowledge graph embedding estimates the $\mathcal{KG}_{\rm OWA}$. Then, the answers are derived by the fuzzy logic inference or optimization. These methods rely on the standard logic calculation and cannot be directly applied to our SQUK setting introduced in Section 3.
3 Soft Queries on Uncertain KG (SQUK)
This section unfolds the concept of soft queries on UKG. We first give the formal definition of UKG and the open world assumption. Then, we present soft queries on UKG, which is motivated by real-world decision-making scenarios in Example 1 and well supported by the long-standing establishment in possibilistic Constraint Satisfaction Problems (CSP) (Schiex, 1992) and soft constraint programming (Bistarelli et al., 1999; Rossi et al., 2006).
3.1 Uncertain knowledge graphs
Uncertain knowledge graphs extend every triple $(s,r,o)\in\mathcal{KG}$ with a confidence value $p\in[0,1]$ (Chen et al., 2019; Pai & Costabello, 2021) [^1] For simplicity, we follow the practice of uncertain knowledge graphs to assume $p\in[0,1]$. The value $p$ can indicate the confidence, importance, or strength of the triple in different domains.. The larger $p$ indicates the higher possibility of the knowledge being true. Formally, an uncertain knowledge graph is defined as follows:
Definition 5 (Uncertain knowledge graph (closed world)).
An uncertain knowledge graph is $\mathcal{UKG}=(\mathcal{KG},P)$. The key to characterize $\mathcal{UKG}$ is the confidence function $P:\mathcal{E}\times\mathcal{R}\times\mathcal{E}\mapsto[0,1]$, which models the confidence value $p=P(s,r,o)$ of the triple $(s,r,o)$. Let $S_{P}=\{(s,r,o):P(s,r,o)>0\}$ be the support set of $P$, we have $S_{P}=\mathcal{KG}$.
The definition of UKG in a closed world assumes that the unobserved knowledge is impossible to be true. The open world assumption for uncertain KG is defined as follows.
Definition 6 (Open world assumption for UKG).
Given $\mathcal{UKG}=(\mathcal{KG},P)$, there is an unobserved $\mathcal{UKG}_{\rm OWA}=(\mathcal{KG},\mathcal{P})$ defined by an open world confidence function $\mathcal{P}:\mathcal{E}\times\mathcal{R}\times\mathcal{E}\mapsto[0,1]$. Let $S_{\mathcal{P}}=\{(s,r,o):\mathcal{P}(s,r,o)>0\}$ be the support set of $\mathcal{P}$, we assume $S_{P}=\mathcal{KG}\subset S_{\mathcal{P}}$.
Remark 1.
The open world assumption does not guarantee $P(s,r,o)=\mathcal{P}(s,r,o)$ when $(s,r,o)\in\mathcal{KG}$, which means that the observed confidence value $P(s,r,o)$ is just an estimation of the underlying uncertainty $\mathcal{P}(s,r,o)$.
3.2 Soft queries: formal definitions
As illustrated in Figure 1, the soft query can be intuitively understood as combining logical structure and soft requirements. In this part, we provide a thorough and formal definition of the entire problem setting of soft query answering.
The soft query is defined as follows.
Definition 7 (Soft EFO query).
A soft EFO query $\Phi$ is
$\Phi=(\Gamma,\Delta_{\Gamma}),$ (5)
where $\Gamma$ is an EFO query defined in Definition 3, and
$\Delta_{\Gamma}=\left\{(\alpha_{ij},\beta_{ij}):\begin{array}[]{c}\alpha_{ij}\in[0,1],\beta_{ij}\in[0,+\infty)\\ i=1,...,q,j=1,...,p_{i}\end{array}\right\}$ (8)
is the set of soft requirements: $\alpha_{ij}$ is the necessity value and $\beta_{ij}$ is the importance value, for the atomic query $a_{ij}$.
Query structure of $\Phi$. The soft query $\Phi$ preserves the query structure of $\Gamma$. It is also syntactically acceptable to expand Equation (5) in the DNF as Equation (1).
@@ 3.2 Soft queries: formal definitions
$\Phi=\phi_{1}\mathchoice{\mathbin{\ooalign{$\displaystyle\lor$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\lor$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\lor$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\lor$\crcr$\scriptscriptstyle\bigcirc$}}}\cdots\mathchoice{\mathbin{\ooalign{$\displaystyle\lor$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\lor$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\lor$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\lor$\crcr$\scriptscriptstyle\bigcirc$}}}\phi_{q},$ (25)
@@
$\phi_{i}=\exists x_{1},...,x_{n}.\psi_{i1}\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\cdots\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\psi_{im_{i}},$ (42)
where $\psi_{ij}=(a_{ij},(\alpha_{ij},\beta_{ij}))$ is the soft atomic query.
Soft entailment under soft requirements. Standard logical calculus and boolean semantics over $\lor$, $\land$, and $\lnot$ are incompatible with the soft query because of the presence of soft requirements in $\Delta$, which means the standard entailment $\models$ does not further apply to soft query answering. Therefore, we define a new operator for soft entailment $\models_{s}$ that evaluates the logical sentences (recall that sentences are queries with all free variables substituted). The output of $\models_{s}$ is not a boolean value but a utility value in $[0,+\infty)$. It means that the query answering problem induced by $\models_{s}$ is to define the answer utility value rather than the previous answer set defined by the boolean semantics of $\models$.
Definition 8 (Answer utility for soft queries).
Given the uncertain KG $\mathcal{UKG}=(\mathcal{KG},P)$ and a soft query $\Phi(y)$, the utility value $U[\Phi,P](s)\in\mathbb{R}_{+}$ of entity $s$ is
$U[\Phi,P](s)=[P\models_{s}\Phi(s/y)].$ (43)
It is also convenient to use a vector $U\in\mathbb{R}_{+}^{|\mathcal{E}|}$ to represent the utility values of all entities, with the $i$-th element the utility value of the $i$-th entity in $\mathcal{E}$.
Then, we introduce the rules of soft entailment operator $\models_{s}$.
Definition 9 (Soft entailment rules).
Given the confidence function $P$, $\models_{s}$ satisfies the following rules:
(1) Soft disjunction.
Let $\phi$ and $\psi$ be soft sentences, $\left[P\models_{s}(\phi\mathchoice{\mathbin{\ooalign{$\displaystyle\lor$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\lor$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\lor$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\lor$\crcr$\scriptscriptstyle\bigcirc$}}}\psi)\right]=\max\left(\left[P\models_{s}\phi\right],\left[P\models_{s}\psi\right]\right)$.
(2) Soft conjunction.
Let $\phi$ and $\psi$ be soft sentences, $\left[P\models_{s}(\phi\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\psi)\right]=\left[P\models_{s}\phi\right]\otimes\left[P\models_{s}\psi\right]$.
(3) Existential variable elimination.
Let $\phi=\exists x.\psi(x)$ where $\psi(x)$ is a soft query with existential variable $x$, $\left[P\models_{s}\phi\right]=\oplus_{s\in\mathcal{E}}\left[P\models_{s}\psi(s/x)\right]$.
(4) Inference soft atomic sentence
Let $\psi=(a,(\alpha,\beta))$ be a soft atomic query with all variables substituted.
(i) positive constraint
$a=r(h,t)$
$\left[P\models_{s}\psi\right]=\beta[P(h,r,t)]_{\alpha};$ (44)
(ii) negative constraint
$a=\lnot r(h,t)$
$\left[P\models_{s}\psi\right]=\beta[1-P(h,r,t)]_{\alpha}.$ (45)
The thresholding operation $[p]_{\alpha}$ is defined as
@@
$\oplus$ and $\otimes$ represent the addition and multiplication in c-semirings. 0 represents the additive identity.

### v2

# Extending Complex Logical Queries on Uncertain Knowledge Graphs

@@ Abstract
The study of machine learning-based logical query answering enables reasoning with large-scale and incomplete knowledge graphs. This paper advances this area of research by addressing the uncertainty inherent in knowledge. While the uncertain nature of knowledge is widely recognized in the real world, it does not align seamlessly with the first-order logic that underpins existing studies. To bridge this gap, we explore the soft queries on uncertain knowledge, inspired by the framework of soft constraint programming. We propose a neural symbolic approach that incorporates both forward inference and backward calibration to answer soft queries on large-scale, incomplete, and uncertain knowledge graphs. Theoretical discussions demonstrate that our method avoids catastrophic cascading errors in the forward inference while maintaining the same complexity as state-of-the-art symbolic methods for complex logical queries. Empirical results validate the superior performance of our backward calibration compared to extended query embedding methods and neural symbolic approaches.
Extending Complex Logical Queries on Uncertain Knowledge Graphs
[…]
@@ 1 Introduction
Representing and reasoning with factual knowledge are essential functionalities of artificial intelligence systems. As a powerful way of knowledge representation, Knowledge Graphs (KGs) (Miller, 1995; Suchanek et al., 2007; Vrandečić and Krötzsch, 2014) use nodes to represent entities and edges to encode the relations between entities. Recently, Complex Query Answering (CQA) over KGs has attracted considerable attention because this task requires multi-hop logical reasoning over KGs and supports many applications (Ren et al., 2023). This task requires answering the existential First Order Logic (FOL) query, involving existential quantification ($\exists$), conjunction ($\land$), disjunction ($\lor$), and negation ($\lnot$). While answering FOL queries has been extensively researched by database community (Riesen et al., 2010; Hartig and Heese, 2007), such studies overlook the incompleteness of most KGs. Consequently, conventional graph traversal methods for relational database queries may neglect certain answers due to the missing links of KGs. In recent studies on complex logical queries on knowledge graphs, the generalizability of machine learning models is leveraged to predict the missing links of observed KGs and conduct first-order logic reasoning (Ren and Leskovec, 2020; Arakelyan et al., 2021; Liu et al., 2021; Wang et al., 2023b). This combination of machine learning and logic enables further possibilities in data management (Ren et al., 2023).
@@
Figure 1: (a) Examples of two soft queries in the candidate search procedure. The soft queries introduced in this paper are jointly defined by first-order logic and soft requirements. In particular, soft requirements (necessity and importance) are introduced to characterize fine-grained decision-making preferences, distinguishing them from first-order queries. (b) Incomplete uncertain KG for to what extent a candidate possesses a skill. Solid lines indicate the observed knowledge, while dashed lines indicate the unobserved data. Values indicate confidence level, where the higher value indicates the fact is more likely to be true.
The uncertainty of knowledge is widely observed ranging from daily events (Zhang et al., 2020) to complex biological systems (Szklarczyk et al., 2023). To represent the uncertain knowledge, confidence values $p$ are associated with facts to augment the KG (Carlson et al., 2010; Speer et al., 2017; Szklarczyk et al., 2023), known as the uncertain KG. As exemplified in the right of Figure 1, which illustrates the Job Candidates uncertain KG, confidence values are used to quantify the degree of practice level associated with specific skills. Uncertainty is prevalent in existing KGs primarily because most are constructed using machine learning models, where the predicted likelihood of relational facts inherently introduces uncertainty. Notable examples include O*NET (Pai and Costabello, 2021), STRING (Szklarczyk et al., 2023), ConceptNet (Speer et al., 2017), and AESR (Zhang et al., 2020). To address the incompleteness of uncertain KGs, recent studies estimate the confidence values of missing facts with the generalization power of ML models (Chen et al., 2019; Pai and Costabello, 2021).
@@
Problem Settings Language of reasoning Uncertainty of Knowledge Unobserved Knowledge
@@
Complex logical queries on KG EFO - ML generalization to unobserved facts
@@
To reason with uncertainty, many extensions of first-order logic have been made to cope with the uncertainty in knowledge representation systems formally Adams (1996). It is noteworthy to mention that the study of probabilistic databases also extends the relational databases with a confidence value $p\in[0,1]$ Cavallo and Pittarelli (1987); Suciu et al. (2022). From a machine learning perspective, however, previous studies in open world probabilistic databases are limited from two aspects: (1) They assume uniform uncertainty for all unobserved knowledge Ceylan et al. (2021), leading to weaker characterization of the incomplete knowledge without the generalization. (2) They focus on first-order logic, which might be insufficient to describe practical reasoning processes with uncertainty. To address the limitations, we extend the complex query answering to uncertain KG and propose soft queries combining query structure and soft requirements, as shown in the left of Figure 1.
This paper studies the machine learning method for reasoning with incomplete and uncertain knowledge, advancing previous studies in symbolic probabilistic databases. Its contribution is threefold.
@@
Contribution 1: A novel and practical setting. We propose a novel setting of Soft Queries on Uncertain KG (SQUK). Our setting extends the previous setting of complex logical queries on KGs in two ways: (1) For the incomplete knowledge base, SQUK extends the incomplete KG to incomplete and uncertain KG. (2) For the language describing reasoning, SQUK extends first-order language to uncertainty-aware soft queries with soft requirements, which are motivated by real-world reasoning with uncertainty and the establishment in soft constraint programming (Schiex, 1992; Rossi et al., 2006). We also introduce the formal definition in Section 3. The comparison of SQUK against other settings is detailed in Table 1.
@@
Contribution 2: ML method for soft queries. We bridge machine learning and SQUK by proposing Soft Reasoning with calibrated Confidence values (SRC), which uses Uncertain Knowledge Graph Embeddings (UKGEs) to tackle the unobserved information and achieves the same computational complexity as the state-of-the-art inference algorithms (Bai et al., 2023; Yin et al., 2024). The error analysis is also conducted for SRC  given the error bound function, characterizing how the performance is affected by UKGEs and the query structures. Based on our analysis, we suggest calibrating the confidence by debiasing and learning, which further boosts the performance of SRC.
@@
Contribution 3: Extensive empirical studies. We also conduct extensive empirical studies to benchmark the performance of a broad spectrum of methods under the UKGE (Chen et al., 2019) settings. The calibrated SRC is compared against baselines including Query Embedding  (Ren and Leskovec, 2020) methods with Number Embeddings (QE+NE) (Vaswani et al., 2017) and symbolic search method (Yin et al., 2024). In particular, we compared the differences between QE+NE and SRC under various soft query settings, demonstrating the advantage of SRC. We also make a fair comparison with large language models on annotated soft queries in a natural language setting.
We highlight the uniqueness of the SQUK setting by examining the differences between uncertain KGs and KGs, comparing soft queries with logical queries, and addressing the challenges posed by two versions of incomplete knowledge. These differences are also summarized in Table 1. Additionally, we present the related work concerning complex logical query answering and uncertain knowledge graph embedding in Appendix A.
2 Background
Uncertain KGs enhance traditional KGs by augmenting each triple fact with a confidence value, thereby facilitating the modeling of uncertain knowledge, which is particularly useful in various domains. Figure 1 illustrates uncertainty in job backgrounds. We formally define an uncertain knowledge graph as a set of knowledge as follows:
Definition 1 (Uncertain knowledge graph).
Let $\mathcal{E}$ be the set of entities and $\mathcal{R}$ be the set of relations, an uncertain knowledge graph $\mathcal{G}$ is a set of quadruple $\{(s_{i},r_{i},o_{i},p_{i})\}$, where $s_{i},o_{i}\in\mathcal{E}$ are entities, $r_{i}\in\mathcal{E}$ is relation and $p_{i}\in[0,1]$ [^1] The values of uncertain KGs also indicate the strength or importance. For simplify, previous work  (Pai and Costabello, 2021) normalized the range of values into the interval $[0,1]$. represents the confidence value for the relation fact $(s_{i},r_{i},o_{i})$. This confidence value $p_{i}$ indicates the degree of certainty regarding the truth of the fact.
Following the closed-world assumption (Reiter, 1981) and treating all unobserved facts as false, we can derive the weight graph form for uncertain KG and represent it with the confidence function $P:\mathcal{E}\times\mathcal{R}\times\mathcal{E}\mapsto[0,1]$ as follows:
$P(h_{i},r_{i},t_{i})=\left\{\begin{array}[]{cc}p_{i}&(h_{i},r_{i},t_{i},p_{i})\in\mathcal{G},\\ 0&{\rm otherwise}.\end{array}\right.$ (1)
Uncertain KGs also suffer incomplete issues (Chen et al., 2019, 2021a), with observed knowledge representing only a small portion of the total facts. The assumption that all unseen relational facts are false is inappropriate in real-world scenarios. To address this challenge, previous research on uncertain KGs has proposed a machine learning task to predict the confidence scores of these unseen relational facts (Chen et al., 2019, 2021a). Typically, the observed knowledge in uncertain KGs is split into three nested sets of facts, where $\mathcal{G}_{\text{train}}\subsetneq\mathcal{G}_{\text{valid}}\subsetneq\mathcal{G}_{\text{test}}$. The training set $\mathcal{G}_{\text{train}}$ is used to train the model, while the validation and test sets are used to evaluate its performance in predicting the confidence scores of unseen facts.
Uncertain Knowledge Graph Embeddings (UKGEs) Chen et al. (2019, 2021a) have been the mainstream methods for predicting unseen relational facts in uncertain KGs, as they learn low-dimensional representations that effectively capture the semantics between relations and facts, demonstrating strong generalizability. UKGEs are trained on partial facts $\mathcal{G}_{\text{train}}$ and approximate the confidence function $\mathcal{P}$ deriving from complete facts, defined as the following confidence function:
Definition 2.
An UKGE parameterizes a differentiable confidence function $\hat{P}:\mathcal{E}\times\mathcal{R}\times\mathcal{E}\mapsto[0,1]$.
In practice, obtaining complete facts is challenging, so $P_{\text{test}}$ induced by $\mathcal{G}_{\text{test}}$ are usually substituted for $\mathcal{P}$, which adhere to previous approaches Chen et al. (2019); Pai and Costabello (2021). We present the connection between this setting and the open-world assumption in Appendix E.
3 Soft Queries
The uncertainty inherent in KGs can be modeled using confidence values for each knowledge. However, current complex logical queries are defined on a boolean basis  [^2] For details on logical queries, please refer to Appendix D., which is not compatible with uncertain KGs. This uncertainty necessitates new definitions for logical operations and answer sets. In this section, we introduce the definition of our extended soft queries.
3.1 Syntax and semantic
Definition 3 (Syntax of soft queries).
Soft queries are the disjunction of soft conjunctive queries $\phi_{i}$:
@@ 3.1 Syntax and semantic
$\Phi(y)=\phi_{1}(y)\mathchoice{\mathbin{\ooalign{$\displaystyle\lor$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\lor$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\lor$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\lor$\crcr$\scriptscriptstyle\bigcirc$}}}\cdots\mathchoice{\mathbin{\ooalign{$\displaystyle\lor$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\lor$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\lor$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\lor$\crcr$\scriptscriptstyle\bigcirc$}}}\phi_{q}(y),$ (18)
where $y$ is the free variable. Each $\phi_{i}(y)$ is the conjunction of the soft atomic formula:
$\phi_{i}(y)=\exists x_{1},\dots,x_{n}.a_{i1}\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}\cdots\mathchoice{\mathbin{\ooalign{$\displaystyle\land$\crcr$\displaystyle\bigcirc$}}}{\mathbin{\ooalign{$\textstyle\land$\crcr$\textstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptstyle\land$\crcr$\scriptstyle\bigcirc$}}}{\mathbin{\ooalign{$\scriptscriptstyle\land$\crcr$\scriptscriptstyle\bigcirc$}}}a_{ij},i=1,...,q,$ (35)
@@
Definition 4 (Substitution).
For a soft query involving variables, the substitution replaces all occurrences of the variable $x$ (or $y$) with any entity $s\in\mathcal{E}$ simultaneously, denoted as $s/x$ (or $s/y$).
We denote $\phi(s)$ for the result of substituting $s$ for the free variable $y$. When all variables in the soft query $\phi$ have been substituted, we refer to it as the substituted query. Next, we define the semantics of the soft queries, starting with the soft atomic formula. Specially, the soft atomic formula involves two novelty concepts: $\alpha$ necessity and $\beta$ importance, which are inspired form soft Constraint Satisfaction Problems (CSPs) Rossi et al. (2006) to manipulate the uncertainty of facts. We introduce the related work of soft CSPs in Appendix F.
1. The $\alpha$ necessity component draws inspiration from possibilistic CSPs (Schiex, 1992) and is designed to capture necessity criteria. It serves the purpose of filtering out unnecessary constraints and involves a thresholding operation. The thresholding operation $[p]_{\alpha}$ is defined as follows:
@@
2. The $\beta$ importance component is influenced by weighted CSPs (Bistarelli et al., 1999) to describe preference, which is the weight employed to adjust the relative significance of different conditions.
Definition 5 (Semantic of soft queries).